How many gallons of a solution of 20% acid must be blended with a solution of 5% acid to obtain 80 gallons of a solution of 12% acid

Found 2 solutions by mananth, jorel1380:
Answer by mananth(16946)

percent ---------------- quantity
Acid I 20 ---------------- x gallons
Acid II 5 ---------------- 80 - x gallons
Total 12.00% ---------------- 80

20x+5(80-x)=12*80
20x+400-5x=960
20x-5x=960-400
15x=560
/15
x=37.33 gallons 20.00%
42.67 gallons 5.00%

Answer by jorel1380(3719)

80 gallons of 12% acid equals 9.6 gallons of pure acid.
x+y=80
.20x+.05y=9.6
Multiply the second equation by 100:
20x+5y=960
x+y=80
-------------
20x+5y=960
5x+5y=400
15x=560
x=37 1/3
y=42 2/3.